A 15.0 L closed vessel contains sulfur trioxide, sulfur dioxide, and oxygen gases that are in equilibrium at 350°C, with Kequal to 1.8x10-5.

2SO3 (g) ⇌ 2SO2 (g) + O2 (g)

If the equilibrium partial pressures of SO2 and O2 are 0.024 atm and 0.057atm, respectively, how many grams of SO3 (MM=80 g/mol) are present in the container? (Hint: Use the ideal gas equation.) 


1
Expert's answer
2021-05-09T00:00:25-0400

2SO3 (g) ⇌ 2SO2 (g) + O2 (g)


Let Partial pressure of "SO_2" is x atm.


So

"k_p=\dfrac{P_{O_2}.P_{so_2}^2}{P_{so_3}^2}=\dfrac{0.057\times 0.024^2}{x^2}\\[9pt]x^2=\dfrac{3.28\times 10^{-5}}{1.8\times 10^{-5}}\\[9pt]\Rightarrow x=\sqrt{1.82}=1.35\text{ atm}"



Now using the ideal gas equation-


"PV=nRT\\[9pt]

\Rightarrow 1.35\times 15=n\times 0.08314\times 623K

\\[9pt]\Rightarrow n=\dfrac{20.25}{51.796}=0.39"


Grams of "SO_3" = moles of "SO_3 \times" Molar mass of "SO_3"

"=0.39\times 80=31.27g"